OpenStudy (anonymous):

Verify the identity. cotangent of x divided by quantity one plus cosecant of x equals quantity cosecant of x minus one divided by cotangent of x Please! Thank you!

OpenStudy (anonymous):

(cosx/ sin x)/ 1+ 1/sin x (cos x/ sin x) * (cos x * sin x/ sin x^2) (cos x * sin x) + (cos x * sin x) /sin^2 x That is as far as I got, but I don't think it is right.

OpenStudy (johnweldon1993):

Well that was getting me nowhere (changing to sin and cos) so I'll just cross multiply and go from there \[\large \frac{cot(x)}{1 + csc(x)} = \frac{csc(x) - 1}{cot(x)}\] Cross multiply \[\large cot^2(x) = (1 + csc(x))(csc(x) - 1)\] Now change everything to sin and cos \[\large (\frac{cos(x)}{sin(x)})^2 = (1 + \frac{1}{sin(x)})(\frac{1}{sin(x)} - 1)\] Put the fractions over common denominators of \(\large sin\) \[\large (\frac{cos(x)}{sin(x)})^2 = (\frac{1 + sin(x)}{sin(x)})(\frac{1 - sin(x)}{sin(x)})\] Multiply that out \[\large \frac{cos^2(x)}{sin^2(x)} = \frac{(1 + sin(x))(1 - sin(x))}{sin^2(x)}\] \[\large \frac{cos^2(x)}{\cancel{sin^2(x)}} = \frac{(1 + sin(x))(1 - sin(x))}{\cancel{sin^2(x)}}\] \[\large cos^2(x) = (1 + sin(x))(1 - sin(x))\] \[\large cos^2(x) = 1 - sin^2(x)\] \[\large cos^2(x) = cos^2(x)\] Woo! :) sorry lot of typing

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (johnweldon1993):

Anytime :)